If sin theta + cos theta = 1 , then find sin 2 theta

Grade:11

1 Answers

Anudh
6 years ago
The max value sin and cos can take is 1.....Therefore it`s like...Sin(90°)+cos(90°)=1+0=0.Therefore the value theta takes is 90°...Hence sin(2theta)=sin(2×90°)=0.The answer is 0.